OpenStudy (anonymous):

If α and B are the measures of two first quadrants angles and sinα=4/5 and sin B= 5/13 find sin (α+B)

OpenStudy (callisto):

sinα = 4/5 => cos α = 3/5 sinB = 5/13 => cosB = 12/13 sin (α+B) = sinαcosB + cosαsinB = (4/5)(12/13) + (3/5)(5/13) = ... ?

OpenStudy (anonymous):

63/65?